# Forecasting from Clinical Textual Time Series: Adaptations of the Encoder and Decoder Language Model Families

### v2

@@ Abstract
Clinical case reports encode rich, temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
**footnotetext: These authors contributed equally to this work
1 Introduction
Clinical risk forecasting is a cornerstone problem in health care decision making that guides eligibility and reasoning about interventions. Because risk prediction performs best with access to rich data streams, great lengths are taken to capture the relevant health factors that affect decision making. However, while many forecasting tasks are performed on structured tabular data, evidence suggests that, despite widespread efforts to collect large amount of data in structured format, only a small fraction of structured data overlaps with that documented in text (Seinen et al., 2025). Furthermore, empirically we observe that predictions are considerably better when text is integrated (Kline et al., 2022).
A complementary data stream in healthcare comes from the textual data that providers use in documenting their encounters. Likely the most comprehensive part are the case reports and discharge summaries, which are a sort of “post-mortem”, retrospective discussion of the relevant factors for decision making and observed progression. However, without further effort, we are unable to leverage this data stream for forecasting because the times that are documented are interwoven in the narrative passage.
To construct and learn from these patient trajectories, this work investigates the development of a forecasting pipeline for textual time series from retrospective analysis of clinical trajectories from case reports. Using the framework in Wang and Weiss (2025), we leverage extracted temporally located clinical findings, namely lists of (event, time) tuples that represent the patient encounter. The mapping from case report to time-localized events allows us to make anytime forecasts by masking those events that occur after the requested forecast time. Without this annotation process, it would be unclear which findings could be used so as to avoid “causal leakage” from events occurring in the future.
Our approach differs from other efforts in the NLP and clinical informatics community and is driven by our particular emphasis on high-performance forecasting and temporally-aware risk characterization. For example, our definition of clinical finding is related to clinical entity and clinical concept definitions from the field of named entity recognition. However, we avoid mapping to a discrete set of entities to increase the availability of contextual information for forecasting, and thus require an alternative annotation guideline (Uzuner et al., 2011). With respect to our temporal annotations, our approach differs from the classical temporal relations task, which, while capturing relative ordering and overlap of events, makes it difficult to perform risk calculations and multimodal alignments that are functions of time (Kohane, 1987; Zhou and Hripcsak, 2007; Sun et al., 2013).
@@ 1 Introduction
We use LLMs to assist us in this task, so it is natural to question whether we need this extraction/annotation step at all, since a multitude of BERT-based and LLM models have consumed these, similar post-mortem reports, and more during training (Peng et al., 2020; Weber et al., 2024). From an architectural perspective, both encoder (BERT-style) and decoder (LLM-style) models will not perform the temporal masking necessary to encode leakage-free embeddings. For BERT models, which use full attention modules with masking, the probability that the random masking recovers a time-ordered representation is vanishingly small (tested empirically on the example timeline in Figure 1 (right) with default BERT masking probability of 0.15; probability of a time-ordered masking: 2/100,000). For decoder models, their causal attention mask means that embeddings are processed in text-order rather than time-order. Given that the concordance, the fraction of the time pairwise temporal ordering are correct, between text-ordered and time-ordered events is lower than 1 in most reports (c=0.23 in the example in Figure 1), we would expect suboptimal temporal reasoning capabilities.
As we will see, we empirically find that these models do not forecast well without further training. If we can address these textual-temporal discrepancies, we expect to be able to substantially improve LLM capabilities at clinical forecasting, which could lead to downstream improvements in clinical temporal reasoning, risk estimation, causal reasoning, and decision making.
1.1 Contributions
  1. 1.
    Annotation and Extraction Pipeline: Building on methods of Noroozizadeh and Weiss (2025) for extracting information from the PubMed Open Access corpus, we develop a pipeline that converts case reports into textual time series using both regex-based filtering and LLM-assisted extraction. The pipeline produces temporally localized (event, time) tuples, mitigating potential causal leakage by ensuring that only pre-cutoff events inform forecasts.
  2. 2.
    Comprehensive Model Comparison: We rigorously evaluate a broad suite of models across multiple forecasting tasks. Our study compares instruction-tuned decoder-based models (both via fine-tuning with MLP heads and prompting strategies) against encoder-based models with task-specific fine-tuning. Experiments on event occurrence prediction, temporal ordering, and survival analysis reveal complementary strengths: no single modeling approach consistently outperforms across all tasks, highlighting the importance of aligning model choice with the specific forecasting objective.
  3. 3.
    Analysis of Temporal versus Text Ordering: We perform a detailed investigation into the impact of ordering on performance by contrasting training on time-ordered versus text-ordered annotations. Our experiments demonstrate that preserving the clinical narrative’s natural order can enhance generalization on external datasets, while explicit time ordering can improve ranking performance under specific conditions.
  4. 4.
    Sensitivity Analysis of Temporal Masking: Through systematic dropout experiments, we analyze the effects of randomly masking portions of the clinical history on forecasting accuracy and event ordering. The findings reveal that while classification performance (F1 score) deteriorates with higher levels of masking, the concordance index remains largely stable—highlighting differing sensitivities of binary prediction and ranking tasks to historical context in our textual time series setup.
Generalizable Insights about Machine Learning in the Context of Healthcare
Extensive research has been in the clinical NLP community towards temporal understanding of clinical narratives (Kohane, 1987; Leeuwenberg and Moens, 2020; Cheng and Weiss, 2023). While much has focused on temporal relations, relative-time event tuples facilitate the use of longitudinal and survival models so long as they can use free-text inputs and outcomes (possibly via mappings to an ontology). Whereas the primary focus in this space is on the entity and temporal extraction tasks, downstream application studies remain infrequent. Our study seeks to close thie gap by applying forecasting methodologies to these data streams to characterize the performance of leading machine learning and NLP approaches. A better empirical understanding of the strength and weakness of these approaches will enable future work to advance improved textual time series methods. Our investigation also highlights that risk estimation and event anticipation is not optimally captured in LLMs (as we demonstrate in contrasts between fine-tuning and prompt-based strategies), suggesting the need for careful investigation of LLMs when applied to assessment of risk.
2 Related Work
@@
Extracting timelines from clinical narratives has been a longstanding challenge in biomedical NLP. Early efforts such as the i2b2 2012 challenge provided annotated datasets for temporal relation extraction, focusing on identifying and ordering events in discharge summaries (Uzuner et al., 2011). Follow-up work constructed timelines from these narratives by linking clinical events to timestamps or temporal expressions (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024). These methods often relied on pre-defined event spans and excerpts, rather than full-document processing. In contrast, following Noroozizadeh and Weiss (2025), we assign timepoints to clinical findings from full-length case reports, allowing for finer temporal resolution. While some works integrate structured EHR data to anchor timelines, our approach focuses on timeline construction and forecasting using text alone, which is essential for corpora such as the PubMed Open-Access Subset, where no structured metadata is available. Our annotation strategy enables direct supervision of event-time alignment, overcoming the limitations of earlier span-only annotations. This aligns with the broader need in sepsis and critical care research to incorporate temporality into clinical phenotyping (Johnson et al., 2018; Kamran et al., 2024; Seinen et al., 2025), especially given the high missingness in structured data (Johnson et al., 2016, 2023).
@@
Incorporating unstructured text (clinical notes, narratives) into predictive models has been explored in various contexts, primarily for outcomes like mortality or readmission. Notably, ClinicalBERT and related transformer models pre-trained on medical notes have achieved state-of-the-art results in predicting hospital readmission and mortality (Huang et al., 2019; Gu et al., 2021). These works highlight that text contains complementary information to structured data, for example, documenting patient symptoms or social factors not in coded form. However, directly modeling text is challenging due to noise and length; researchers often rely on the full text or bag-of-words representations, which lose temporal granularity. Our approach differs by preserving the chronological sequence of events mentioned in text, effectively treating narrative data as a time series. This connects to recent studies on re-purposing LLMs for traditional time-series forecasting tasks: techniques like Time-LLM have shown that language models can be “re-programmed” to handle numeric sequences by mapping time points to token sequences (Jin et al., 2023). In the clinical domain, our work is the first to apply language models to text-derived time series, bridging free-text analysis with time-aware prediction. We also relate to prior sepsis prediction models. Traditional models use vital sign trends or clinical scores (e.g., SOFA, SAPS) to predict sepsis onset or mortality (Hou et al., 2020; Noroozizadeh et al., 2023). Unlike these, our method relies solely on narrative descriptions rather than real-time vitals, and is thus complementary: it demonstrates how knowledge embedded in case reports (which often include the clinicians’ interpretations and context) can be harnessed for prediction. This could be especially useful in settings where detailed narratives are available (such as case report literature or perhaps patient histories) but structured data is sparse.
@@
The advent of powerful LLMs like GPT-3 and GPT-4 has sparked interest in various healthcare applications, from medical question-answering to clinical note summarization. These models, trained on massive text corpora, contain a wealth of medical knowledge and the ability to perform reasoning to some extent. For instance, GPT-4 has been tested on medical board exam questions with impressive results, and pilot studies show LLMs can abstract patient records for quality reporting (Boussina et al., 2024). Nevertheless, using LLMs for forecasting in healthcare is relatively uncharted territory. A limitation of out-of-the-box LLMs is their tendency to generate plausible-sounding but unverified content (hallucinations) and their lack of alignment with specific clinical data distributions. Our benchmark explicitly evaluates LLMs through prompting (zero-shot and few-shot) to see if they can ”forecast” events when given a patient’s story so far. We also compare that against fine-tuned models that are directly trained on the task, to quantify the gap between general-purpose and task-specific modeling in this context. This comparison adds to the growing body of evidence that, while large models have broad knowledge, domain-specific fine-tuning or structuring of input is often necessary for reliable performance in clinical tasks (Huang et al., 2019; Gu et al., 2021). Some recent analyses suggest that medical-domain fine-tuning may not consistently improve performance over foundation models (Jeong et al., 2024), further motivating the need for task- and input-specific evaluation setups like ours.
3 Methods
 [^1] We will be releasing our code upon publication.
We outline the task specification, define key terms, describe the extraction and annotation process, and present the evaluation criteria (Figure 2).
Figure 2: Flowchart of the T2S2 forecasting analysis pipeline
3.1 Dataset extraction
@@
We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Building on the method described in Wang and Weiss (2025), we extract the body of each case by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus. To maintain relevance, we include only those documents that contain case-insensitive matches to the regular expressions (case report|case presenta) and year-? ?old. Prior work has shown that this regex-based approach demonstrated notably higher specificity compared to PubMed metadata on a random sample of 20 case reports Noroozizadeh and Weiss (2025).
@@
As demonstrated in prior work Noroozizadeh and Weiss (2025), we applied additional filtering to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
@@
This multi-step process yielded a total of 2319 Sepsis-3 case reports. From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) used for broader testing (Figure 2).
3.2 Textual time-series annotation
@@
As established in previous work Noroozizadeh and Weiss (2025), we aimed to construct a textual time series case report corpus of Sepsis-3 patients from the PMOA sepsis case reports. A textual time series refers to a sequence of clinical findings, each paired with a timestamp (either absolute or relative to the case presentation time), corresponding to an individual patient. Here, a clinical finding is defined as a free-text expression describing an entity that pertains to or may impact a person’s health.
@@
A textual time-series is different from conventional approaches to clinical concept annotation (e.g., Sun et al. (2013); Uzuner et al. (2011)) by extending the span of extracted text to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, coordinated phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
@@
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepis (T2S2) The exact LLM extraction prompt can be found in Appendix E. Our dataset was split into training (n = 1795), testing T2S2-test (n = 244) and the 2 external validation sets sepsis 10 (n = 10) and sepsis 100 (n = 90). The training set was further split into 80% for model training and 20% for model validation (e.g. early stopping).
3.3 Task Definition
@@
Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task:
Event Occurrence Prediction:
@@
Temporal Ordering Prediction:
@@
Survival Analysis for Mortality Time:
@@
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., “the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate performance using the time-dependent concordance index to assess how well the predicted survival times align with actual outcomes.
3.4 Modeling Approaches
@@
To evaluate performance on the event forecasting and survival prediction tasks, we implement three modeling paradigms: (i) large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), and (iii) BERT-style encoder models with task-specific fine-tuned heads. In Appendix C and  D, we provide detailed explanations of how we approach the event occurrence prediction, temporal ordering prediction, and survival analysis tasks using encoder-only models.
Fine-tuned LLMs:
@@
We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants. Each model is paired with a lightweight multilayer perceptron (MLP) head trained for classification or ranking, depending on the task. The input consists of a text-formatted prefix of events (e.g., a clinical timeline up to time $t$), optionally accompanied by an instruction template. The output layer produces task-specific predictions: binary labels for event occurrence or a permutation over $k$ events for ordering. Training is conducted using cross-entropy loss for classification and pairwise ranking loss for ordering.
Prompted LLMs:
@@
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix F. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
Fine-tuned Encoder-Only Models:
@@
We also evaluate a range of encoder-based models trained end-to-end on each task. Architectures include BERT-base-uncased, RoBERTa-base, DeBERTa-v3-small, ModernBERT-small, and ModernBERT-large. For each model, we append a task-specific MLP head for event occurrence or ordering prediction. The model input is the same flattened prefix of the event sequence used in other settings, tokenized and formatted according to the respective architecture. These models are trained using standard supervised learning objectives and evaluated on the same metrics as other methods: F1 score for event occurrence, and pairwise concordance for temporal ordering. For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index.
4 Results

### v3

# Forecasting Clinical Risk from Textual Time Series: Structuring Narratives for Temporal AI in Healthcare

@@ Abstract
Clinical case reports encode temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
Introduction
Healthcare disparities persist globally, with preventable deaths from conditions like sepsis disproportionately affecting underserved populations who often present to under-resourced facilities with limited access to specialized expertise. In these settings, much of the critical diagnostic and prognostic information exists only in unstructured clinical narratives—case reports, discharge summaries, and progress notes—as comprehensive structured data infrastructure may be unavailable (Anzalone et al. 2025; Seinen et al. 2025). While machine learning approaches have shown promise on structured data, with recent work demonstrating that incorporating text alongside structured inputs significantly improves predictive performance (Kline et al. 2022), large language models struggle with clinical risk estimation, especially in patient-facing scenarios (Wong et al. 2025). This gap highlights the need for automated risk forecasting from textual clinical records that could democratize access to expert-level clinical reasoning, particularly in resource-constrained environments where timely specialist consultation is not available. The challenge of LLM risk estimation motivates a structured treatment of risk forecasting and survival modeling from unstructured narrative sources.
Among textual sources, retrospective case reports serve as a rich training ground for developing temporally-aware AI systems that can eventually be deployed on real-time clinical notes. Case reports provide holistic accounts of clinical trajectories with explicit temporal reasoning—exactly the type of structured thinking needed for real-time clinical decision support. However, their narrative structure interleaves temporally unordered information, making them difficult to apply in forecasting tasks. Without identifying the time of occurrence for each event, models are prone to causal leakage—using information not available at the time of prediction. Solving this temporal reasoning challenge in case reports establishes the foundation for deployment in live clinical environments where such capabilities could improve patient outcomes.
@@ Introduction
One might ask whether large pretrained language models can already perform forecasting from clinical narratives, given their exposure to biomedical texts (Peng, Chen, and Lu 2020; Weber et al. 2024). However, these models have architectural limitations for temporal reasoning tasks. Encoder-based models apply random masking that rarely yields temporally coherent representations, while decoder-style models mask text in linear sequence, modeling events in text order rather than time order. Our experiments demonstrate these limitations quantitatively–for example, using the timeline in Figure 1, the probability of recovering a valid time-ordered mask is $0.002\%$, and decoder models achieve low temporal concordance (e.g., $c=0.23$), leading to suboptimal forecasting accuracy without task-specific restructuring.
To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples and makes explicit the temporal structure of patient trajectories. Building on recent annotation work (Wang and Weiss 2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
Our work builds on extensive research in clinical NLP and temporal reasoning (Kohane 1987; Leeuwenberg and Moens 2020; Cheng and Weiss 2023), but shifts the emphasis from extraction to downstream clinical forecasting. While prior efforts have focused on entity recognition and temporal relation extraction (Zhou and Hripcsak 2007; Uzuner et al. 2011; Sun, Rumshisky, and Uzuner 2013), few have systematically assessed how such representations influence downstream tasks such as risk prediction and survival modeling. We bridge this gap by formalizing textual time series representations and evaluating their effectiveness for temporally grounded clinical prediction.
We empirically assess this framework across multiple tasks and modeling strategies, benchmarking encoder and decoder models—both fine-tuned and prompted versions—on three forecasting tasks: event prediction, survival analysis, and temporal ordering of future clinical events. Our results show no single model dominates across all tasks, highlighting the importance of aligning model architecture with forecasting objectives. We further examine how input ordering (time vs. narrative) impacts generalization and analyze the sensitivity of models to missing historical context through systematic masking experiments.
In summary, this paper makes the following contributions: (i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks. Results show no single model excels universally, underscoring the need to align model choice with the specific forecasting objective; (iii) Temporal- versus Text- Ordering: We investigate the impact of annotation order by comparing training on time-ordered versus text-ordered data. Results show that preserving the narrative’s natural order can improve generalization to external datasets, while time-based ordering can enhance ranking performance; (iv) Sensitivity Analysis of Temporal Masking: We conduct systematic dropout experiments by randomly masking parts of the clinical history to assess their impact on forecasting and event ordering. While higher masking levels reduce classification performance (F1), the concordance index remains largely stable, highlighting differing sensitivities of binary prediction and ranking tasks to historical context in textual time series; (v) Methodological Framework for Temporal Clinical Forecasting: Our approach demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks, providing a replicable methodology that could be adapted to other clinical and non-clinical text sources.
Related Works
@@
Extracting timelines from clinical narratives is a challenging biomedical NLP task. The i2b2 2012 challenge introduced annotations datasets for temporal relation extraction from discharge summaries (Uzuner et al. 2011). Subsequent methods linked clinical events to timestamps or temporal expressions (Leeuwenberg and Moens 2020; Frattallone-Llado et al. 2024), typically using pre-defined event spans. We adopt the approach of Noroozizadeh and Weiss (2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution. Unlike methods using structured EHR data, we focus solely on text—crucial for sources like PubMed that lack structured metadata. By directly supervising event-time alignment, we overcome limitations of span-based annotations (Rosenbloom et al. 2011). This aligns with growing emphasis on temporality in sepsis and critical care phenotyping (Johnson et al. 2018; Kamran et al. 2024), particularly given high missingness rates in structured data (Johnson et al. 2023; Seinen et al. 2025).
@@
Clinical text has been used for outcome prediction tasks like mortality and readmission, with models like ClinicalBERT achieving high levels of performance on EHR notes (Huang, Altosaar, and Ranganath 2019; Gu et al. 2021). Text captures complementary information—symptoms and social factors—not found in structured codes. However, traditional approaches using full-text or bag-of-words representations obscure temporal dynamics. Our approach preserves event order within narratives, treating data as time series. This aligns with recent efforts adapting LLMs for time series tasks, such as Time-LLM mapping numeric sequences to tokens (Jin et al. 2023). To our knowledge, this is the first work applying LLMs to text-derived time series in clinical domains. While prior sepsis prediction models rely on vitals or scores like SOFA and SAPS (Hou et al. 2020; Noroozizadeh, Weiss, and Chen 2023), our method is complementary—leveraging narrative descriptions that include clinician interpretations and context. This is valuable when case reports or patient histories are available but structured data is sparse or unavailable.
@@
Recent LLMs such as GPT-3 and GPT-4 have enabled applications from medical QA to note summarization. These models encode substantial medical knowledge and reasoning capacity—GPT-4 demonstrates strong performance on board exams and can support patient record abstraction for quality reporting (Boussina et al. 2024). However, their use in clinical forecasting remains largely unexplored. Our benchmark tests LLMs via zero- and few-shot prompting to assess clinical event forecasting from patient narratives, comparing results to fine-tuned models to quantify the gap between general-purpose and task-specific modeling. This contributes evidence that while LLMs hold broad medical knowledge, fine-tuning and structured inputs are often required for clinical reliability (Huang, Altosaar, and Ranganath 2019; Gu et al. 2021). Some studies suggest domain-specific tuning does not always improve over foundation models (Jeong et al. 2024), reinforcing the need for task-specific evaluations like ours.
Methods [^1] Our code is available at https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting
We next describe our dataset construction and annotation pipeline, define the forecasting and survival prediction tasks, detail the modeling approaches evaluated, and present the sensitivity analyses used to assess robustness (Figure 2).
Figure 2: T2S2 forecasting analysis pipeline
Dataset Extraction
@@
We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Building on the method described in prior work (Wang and Weiss 2025), the body of each case was extracted by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus. To maintain relevance, we include only documents containing case-insensitive matches to (case report|case presenta) and year-? ?old, an approach shown to have higher specificity than PubMed metadata (Noroozizadeh and Weiss 2025).
@@
As demonstrated in prior work of (Noroozizadeh and Weiss 2025), additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
@@
This multi-step process yielded a total of 2319 Sepsis-3 case reports. From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) as a bronze standard for broader testing based on an alternative annotator (O1-preview; Figure 2).
Textual Time Series Annotation
@@
Following Noroozizadeh and Weiss (2025), a textual time series case report corpus of Sepsis-3 patients was constructed from the PMOA sepsis reports. A textual time series refers to a sequence of clinical findings, each paired with a timestamp (either absolute or relative to the case presentation time), corresponding to an individual patient. Here, a clinical finding is defined as a free-text expression describing an entity that pertains to or may impact a person’s health.
@@
A textual time series differs from conventional approaches to clinical concept annotation (e.g., (Sun, Rumshisky, and Uzuner 2013; Uzuner et al. 2011)) in its extension of the text span to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, coordinated phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
@@
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepsis (T2S2). The exact LLM extraction prompt can be found in Appendix E. Our dataset was split into training (n = 1,795; 80% train, 20% validation), testing T2S2-test (n = 244), and two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).
Forecasting Tasks
@@
Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task as follows:
Event Occurrence Prediction.
@@
Temporal Ordering Prediction.
@@
Survival Analysis for Mortality Time.
@@
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., ”the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate using the time-dependent concordance index to measure alignment of the predicted survival times with actual outcomes.
Modeling Approaches
@@
To evaluate performance on the event forecasting and survival prediction tasks, we implement five modeling paradigms: (i) decoder-only large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), (iii) encoder-only models with task-specific fine-tuned heads, (iv) encoder-masking models with fine-tuning, and (v) encoder-masking models in zero-shot settings. In Appendix C and  D, we provide detailed explanations of implementation details.
(i) Fine-tuned LLMs.
We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants (Grattafiori et al. 2024; Liu et al. 2024). We also include open-source models with documented training corpora (OLMO-32B-Instruct (OLMo et al. 2024), RedPajama-INCITE-7B-Instruct (Weber et al. 2024)) to address potential data leakage concerns–notably, RedPajama explicitly excludes PubMed from its training data. Additionally, we evaluate a medically fine-tuned decoder model (MediPhi-PubMed (Corbeil et al. 2025)) to assess domain-specific benefits.
@@
Each model is paired with a lightweight multilayer perceptron (MLP) head trained for classification or ranking, depending on the task. The input consists of a text-formatted prefix of events (e.g., a clinical timeline up to time $t$), optionally accompanied by an instruction template. The output layer produces task-specific predictions: binary labels for event occurrence or a permutation over $k$ events for ordering. Training is conducted using cross-entropy loss for classification and pairwise ranking loss for ordering.
(ii) Prompted LLMs.
@@
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix E. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
(iii) Fine-tuned Encoder-Only Models.
@@
We also evaluate a range of encoder-based models trained end-to-end on each task. Architectures include general-purpose models (BERT-base-uncased (Devlin et al. 2019), RoBERTa-base (Liu et al. 2019), DeBERTa-v3-small (He, Gao, and Chen 2021), ModernBERT-base/large (Warner et al. 2024)) and biomedically-pretrained variants (BioClinical-ModernBERT-base/large (Sounack et al. 2025)). For each model, we append a task-specific MLP head for event occurrence or ordering prediction. The model input is the same flattened prefix of the event sequence used in other settings, tokenized and formatted according to the respective architecture. These models are trained using standard supervised learning objectives and evaluated on the same metrics as other methods: F1 score for event occurrence, and pairwise concordance for temporal ordering. For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index (Antolini, Boracchi, and Biganzoli 2005).
(iv, v) Encoder-Masking Models.
We adapt masked language modeling (MLM) for our temporal reasoning tasks. For event occurrence, models predict masked tokens (“yes”/“no”) in prompts such as “Will [event] happen within 24 hours? [MASK]”. For temporal ordering, models predict “before”/“after” tokens in prompts comparing event pairs. We evaluate both fine-tuned variants (trained on task-specific objectives) and zero-shot variants (using pretrained MLM capabilities without additional training). This approach leverages the bidirectional context of encoder models while maintaining interpretable predictions through constrained vocabulary.
Sensitivity Analyses
We conduct sensitivity analyses to examine how the temporal structure and completeness of input sequences affect forecasting performance. These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information:
Time-ordering Strategies.
We evaluate two strategies for ordering clinical events within textual time series inputs. In the text-ordered setting, events are presented in the order they appear in the original case report narratives, preserving the narrative structure produced during extraction. In the time-ordered setting, events are sorted chronologically by their documented occurrence time, enforcing strict temporal alignment. This comparison isolates the effect of narrative sequencing versus explicit temporal ordering on model performance. For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder-based models.
Timestep Dropout.
To assess robustness to incomplete patient histories, we introduce a timestep dropout rate (TDR), defined as the proportion of input timesteps randomly removed prior to inference. This procedure simulates varying degrees of missing clinical documentation that may arise in deployment settings. We vary TDR from 0% (full history) to 90% (severely truncated history) in increments, masking events independently and uniformly at random. This setup enables controlled evaluation of model performance under partial information, probing the extent to which predictive accuracy and event ordering depend on the completeness of historical context.
Together, these analyses characterize model robustness to disruptions in temporal coherence and input completeness – two key challenges in real-world deployment scenarios.